If 4x – 7 = 11, then x = 4.
If 4x – 7 = 11, then x = 4.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The Given statement is False

Consider the equation, 4x–7 = 11
Transpose -7 to RHS then it becomes 7.
4x = 11 + 7 4x = 18
x = 18/4
x = 9/2

Thanks